# Driving with Prior Maps: Unified Vector Prior Encoding for Autonomous Vehicle Mapping

### v3

@@ 4.1 Experimental settings
Furthermore, to evaluate the effectiveness of our method across different evaluation metrics and for the rasterized results, we follow the previous works [15, 35, 13] and employ IoU as the metrics for segmentation results.
@@
For fair comparisons, experiments were conducted on all baseline models using their original configs and parameters without any changes. All experiments were trained with 6 NVIDIA RTX A6000. For pre-training UVE, we trained the 24-epoch on nuScenes vecror data. For historical prediction maps, we run each base model to get their respective online local maps.
@@
MGMap [22] 52.8 67.5 68.1 62.8
MapQR [26] 60.1 71.2 66.2 65.9
MapTRv2 [20] 60.7 68.9 64.5 64.7
# SD Prior (Ours) 62.0 70.2 66.3 66.2
# Online Local Prior (Ours) 63.9 71.7 68.6 68.1
# HD map-EX* (Ours) 61.5 69.8 69.4 66.9
Table 2: Comparison on Argoverse 2 where the 3D map elements are predicted directly.
4.2 Main results
@@
To evaluate the effectiveness of our method across different model, evaluation metrics, and vector prior maps, we choose HDMapNet and MapTRv2 as baselines (see Table 1). SD maps, which primarily consists of a centerline skeleton, showed enhancements of 3.0 mAP for MapTRv2 and 2.3 mIoU for HDMapNet, respectively. And for query-based model, our method has a greater improvement because our HPQuery makes better use of prior maps. Online local maps, which offer up-to-date information, led to more significant improvements, with 4.2 mAP and 2.6 mIoU for MapTRv2 and HDMapNet, respectively. These results indicate that more accurate and current local maps contribute to better performance by providing more comprehensive and precise information. Figure 5 illustrates how online local maps effectively restore obscured map elements. More visualization are shown in the supplementary material. Although offline HD maps are highly accurate, their lack of complete prior information resulted in notable improvements of 3.6 mAP and 3.2 mIoU for MapTRv2 and HDMapNet, respectively. Additionally, we evaluated the results using a stricter threshold of 0.5m, achieving improvements of 4.1 mAP, 6.1 mAP, and 5.7 mAP, respectively, underscoring that our method provides more substantial improvements under stricter evaluation conditions. Overall, our method demonstrated consistent performance improvements across all baselines with various prior maps, highlighting its generalizability and potential applicability to other mapping frameworks.
Method $\textbf{DET}_{l}$ $\textbf{DET}_{t}$ $\textbf{TOP}_{ll}$ $\textbf{TOP}_{lt}$ OLS
MapTR [19] 17.7 43.5 5.9 15.1 31.0
TopoNet [17] 28.6 48.6 10.9 23.8 39.8
TopoLogic [7] 29.9 47.2 23.9 25.4 44.1
# SD Prior (Ours) 29.9 47.1 29.6 32.9 47.2
# Online Local Prior (Ours) 31.0 47.8 30.9 32.5 47.9
Table 3: Performance comparison with SOTA methods on OpenLaneV2 subset_A set, where SD Prior comes from the SD map of the openLanev2 itself. Since the map elements here have only centerline and no boundaries, there is no HD map-EX* prior.
@@
In Table 2, we present the results of Argoverse 2, which provides 3D vectorized map elements with additional height information as ground truth. These testings are all trained 6 epochs with ResNet50 as the backbone. Our PriorDrive improves performance by 1.5, 3.4, and 2.2, respectively, using different prior maps, demonstrating the superior generalization ability of PriorDrive in terms of 3D vectorized map construction.
@@
We compared PriorDrive with SOTA methods when using ResNet-50 for training 24 epochs. Table 3 shows the results on subset A of OpenLane-V2. Compared to TopoLogic, our method achieves better topological reasoning accuracy using connected SD priors (29.6 v.s. 23.9 on $TOP_{ll}$, 32.9 v.s. 25.4 on $TOP_{lt}$), and the overall score OLS increased by 3.1, indicating that connected SD is more beneficial for lane connectivity and correspondence with traffic elements. Employing online local priors resulted in decent detection accuracy for both detection and topological reasoning, with OLS rising to 47.9. Overall, PriorDrive is a plug-and-play solution that can be widely applied to other vector tasks.
M N Dim IoU
@@
4 4 64 20.7 41.3 41.1 34.4
6 2 64 20.5 41.1 40.9 34.2
2 6 64 20.7 41.0 41.3 34.3
2 2 128 19.5 41.4 41.2 34.0
2 2 64 20.8 41.7 41.8 34.8
1 1 32 19.8 41.1 40.9 33.9
Table 4: Ablation experiments on UVE structure within HDMapNet and SD map prior.
Method Dist Err AP
@@
Without Pretrain - 61.3 63.4 64.4 63.0
Original - - 0.09m 61.2 64.1 64.5 63.3
Noise 3m 20% - 0.46m 61.4 63.0 63.8 62.7
Noise 1m 10% - 0.15m 61.6 63.8 64.2 63.2
Noise 1m 10% 5% 0.18m 61.6 64.3 64.8 63.6
Noise 1m 10% 5% 0.20m 62.5 66.3 64.5 64.4
Mask 10% 5% 0.43m 61.8 64.0 64.3 63.4
Table 5: Ablation of the pre-training strategies. Here, we fixed prior_num=multiple, search_range=5m and integration=replace. “Seg” and “Pt” represent segment-level and point-level, respectively. * indicates data augmentation to expand the data volume.
4.3 Ablation study
In this section, we do extensive research on UVE structure, pre-training, number and search range of prior maps, and integration methods. we conduct experiments using online local map priors on MapTRv2 unless otherwise specified.
Structure of UVE.
Table 4 presents the results of ablation experiments on the UVE structure using the SD prior map within HDMapNet. We observed that increasing the number of layers or expanding the feature dimension of UVE yielded only marginal improvements, averaging around 1.3 mIoU. This limited gain may be attributed to the model’s excess capacity, leading to overfitting. Conversely, a minimal configuration of UVE, with only one layer each of intra-vector and inter-vector attention and a dimension of 32, resulted in a modest improvement of just 1 mIoU. However, when UVE was configured with two layers of intra-vector and inter-vector attention and a feature dimension of 64, the model’s performance reached saturation.
@@ 4.3 Ablation study
- None 59.8 62.4 62.4 61.5
search_range=5m One 61.1 63.8 63.6 62.8
pretrain=true Two 61.4 66.2 63.0 63.5
integration=replace Multiple 62.5 66.3 64.5 64.4
Table 6: Ablation experiments on the number of online local prior maps on MapTRv2.
@@
- None 59.8 62.4 62.4 61.5
prior_num=2 5 m 62.5 66.3 64.5 64.4
pretrain=true 10 m 61.4 64.6 64.4 63.5
integration=replace 15 m 61.9 63.2 63.9 63.0
@@
The effect of pre-training.
To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies in Table 5. When we fed the original vector map directly to UVE without any noise, UVE converged with a minimal error of 0.09 meters. However, the simple task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.3 mAP improvement. We added segment-level Gaussian noise with a 3-meter standard deviation to 20% of the map elements randomly. The UVE’s average Euclidean distance error was 0.46 meters. However, this approach led to a slight decrease in mAP, likely because the UVE had not fully converged. Next, we applied 1-meter segment-level noise to 10% of the map elements randomly, resulting in a reduced error of 0.15 meters for UVE and a 0.2 mAP improvement compared to the model without pre-training. Building on this, we introduced 1-meter point-level noise to 5% of the vector points. Although this increased UVE’s error to 0.18 meters, it yielded a more substantial mAP improvement of 0.6. This demonstrates that pre-training strategies incorporating both segment-level and point-level noise are more effective at learning the prior information of vector maps. We also experimented with masking instead of noise, which increased UVE’s error to 0.43 meters but still resulted in a 0.4 mAP improvement. Although various proposed tasks can effectively pre-train UVE, they are still limited by the amount of pre-training data. Therefore, we employ horizontal and vertical flipping data augmentation to expand the pre-training dataset. The result of the data augmentation is shown as *, effectively improving 1.4 mAP.
@@
- None 59.8 62.4 62.4 61.5
prior_num=2 Replace 62.5 66.3 64.5 64.4
pretrain=true Add 63.3 64.1 64.4 63.9
search_range=5m Concat 63.6 67.3 66.2 65.7
@@
Impact of the number of prior maps.
We conducted an ablation study to assess the effect of the number of prior maps on model performance (see Table 6). The results indicate that using a single prior map led to a 1.3 mAP improvement, demonstrating its effectiveness in enhancing current perception. Incorporating two prior maps resulted in a more substantial improvement of 2.0 mAP, reflecting the value of richer prior information. Finally, multiple prior maps effectively improved 2.9 mAP, indicating that more prior maps can provide more comprehensive necessary information.
Effect of different search ranges for prior maps.
We performed ablation experiments to explore the impact of search ranges for prior maps in Table 7. The results indicate that increasing the search range leads to a slight decrease in improvement. This suggests that while a larger search range allows the prior map to cover a wider area, it also introduces greater positional error. On the other hand, a more precise, narrower search range may not always provide multiple prior maps for most locations, limiting the benefits.
@@
To evaluate the versatility of our approach, we tested various integration methods in Table 8. Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 2.9 and 2.4, respectively, but these methods may impact the performance of the original query. In contrast, the concatenation method achieved a more substantial improvement of 4.2 mAP. This approach effectively preserves the original query’s integrity while also leveraging the instance-level and point-level prior features learned by UVE, making it the most effective integration method.
Inference speed, memory, model size, and detailed runtime.
Table 9 compares our method with MapTRv2 in the above aspects. The results show that our method maintains similar FPS, GPU memory usage, and parameter count compared to MapTRv2, while achieving a 4.2 mAP improvement. Additionally, we provide the detailed runtime of each component. Our proposed UVE accounts for only 15.3% of the total pipeline runtime, with the majority of inference time spent on the 2D Backbone (50.2%), highlighting the efficiency and practical value of our approach across various application scenarios.
@@
Table 9: Detailed comparison of our method with MapTRv2. FPS and detailed runtime of each component are measured on one A6000 GPU with batch size as 1.
Figure 5: Qualitative results with and w/o online local prior.
5 Conclusion
In this paper, we introduced PriorDrive, a novel framework that effectively leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction. Central to PriorDrive is UVE, which can encode diverse vector data efficiently. Through comprehensive experiments, we demonstrated that UVE, combined with proposed pre-training strategies, significantly improves performance of state-of-the-art models in various vector tasks. PriorDrive not only addresses the challenges associated with online HD map construction in complex environments but also offers a scalable solution that continually improves accuracy over time. The iterative use of historical prediction maps as priors leads to progressively refined map outputs, providing latest road information for autonomous driving.

### v4

@@ Introduction
Method Prior Map Backbone Epoch Metric Ped. Div. Bou. Mean
HDMapNet [ICRA21] (Li2021HDMapNetAO) $\times$ Effi-B0 30 IoU 18.7 40.6 39.5 32.9
P-MapNet [RAL24] (Jiang2024PMapNetFM) $\checkmark$ Effi-B0 30 IoU 22.6 44.1 43.8 36.8
NMP [CVPR23] (Xiong2023NeuralMP) $\checkmark$ Effi-B0 30 IoU 21.0 44.2 46.1 37.1
PriorDrive (Ours) $\checkmark$ Effi-B0 30 IoU 25.1 46.7 48.4 40.1
MapTR [ICLR23] (liao2022maptr) $\times$ R50 24 AP 41.2 49.5 51.1 47.3
P-MapNet [RAL24] (Jiang2024PMapNetFM) $\checkmark$ R50 24 AP 43.7 50.9 53.5 49.4
StreamMapNet [WACV24] (Yuan_2024_streammapnet) Temporal R50 24 AP 60.4 61.9 58.9 60.4
PriorDrive (Ours) $\checkmark$ R50 24 AP 60.9 65.3 63.8 63.3
MapTRv2 [IJCV24] (maptrv2) $\times$ R50 24 AP 59.8 62.4 62.4 61.5
PrevPredMap [WACV25] (peng2024prevpredmap) $\checkmark$ R50 24 AP 64.5 66.9 67.6 66.3
HRMapNet [ECCV24] (zhang2024hrmapnet) $\checkmark$ R50 24 AP 65.8 67.4 68.5 67.2
HRMapNet [ECCV24] (zhang2024hrmapnet) $\checkmark$ R50 110 AP 72.0 72.9 75.8 73.6
InteractionMap [CVPR25] (wu2025interactionmap) Temporal R50 24 AP 69.7 72.7 73.0 71.8
InteractionMap [CVPR25] (wu2025interactionmap) Temporal R50 110 AP 75.3 75.6 77.0 76.0
PriorDrive (Ours) $\checkmark$ R50 24 AP 71.8 81.4 74.1 75.8
PriorDrive (Ours) $\checkmark$ R50 110 AP 78.0 82.0 81.6 80.5
Table 1: Performance of various online mapping models with unified prior maps on nuScenes. All methods use camera input. In each block, the first row represents the baseline model, and all other methods are modifications upon the baseline. “Temporal” means using temporal information. The results of StreamMapNet and HRMapNet are from HRMapNet, while the other results are from the respective papers.
Method Epoch $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
MapTRv2 [IJCV24] 6 60.7 68.9 64.5 64.7
HRMapNet [ECCV24] 30 65.1 71.4 68.6 68.3
PrevPredMap [WACV25] 6 64.1 71.4 67.4 67.6
InteractionMap [CVPR25] 6 66.6 75.6 72.7 71.6
PriorDrive (Ours) 6 68.3 75.8 74.2 72.8
Table 2: Performance comparison of 3D map elements on Argoverse 2.
Method Prior Map $\textbf{DET}_{l}$ $\textbf{DET}_{t}$ v1.0.0 v2.1.0
$\textbf{TOP}_{ll}$ $\textbf{TOP}_{lt}$ OLS $\textbf{TOP}_{ll}$ $\textbf{TOP}_{lt}$ OLS
TopoNet + OSMR [IROS24] (osmr) $\checkmark$ 30.6 44.6 7.7 22.9 37.7 - - -
RoadPainter [ECCV24] (Ma2024RoadPainterPA) $\checkmark$ 36.9 47.1 12.7 25.8 42.6 - - -
SMERF [ICRA24] (SMERF) $\checkmark$ 33.4 48.6 7.5 23.4 39.4 15.4 25.4 42.9
TopoLogic [NeurIPS24] (fu2024topologic) $\times$ 29.9 47.2 18.6 21.5 41.6 23.9 25.4 44.1
PriorDrive (Ours) $\checkmark$ 31.5 48.4 24.6 30.7 46.2 31.4 33.7 48.5
Table 3: Performance comparison with SOTA non-prior/prior-based methods on OpenLane-V2 subset_A set. ’-’ denotes the absence of relevant data.
For the rasterized results, we follow the previous works (Xiong2023NeuralMP; 202506.2087; chen2025visrl; chen2025sifthinker) and employ IoU as the metric for segmentation results.
@@
For fair comparisons, experiments were conducted on all baseline models using their original hyperparameters. All experiments were trained with 8 NVIDIA RTX A6000. For pre-training UVE, we trained the 24-epoch on nuScenes vector data. This one-time offline pre-training ( 12h) yields significant improvements, and its cost is amortized through plug-and-play reuse across various models. The sources of the three types of prior maps are as follows: The SD map is from OpenStreetMap (Haklay2008OpenStreetMapUS); as for the existing HD prior maps, we follow MapEX (MAPEX). For each map element localization, we add noise from a Gaussian distribution with a standard deviation of 1 meter. This has the effect of applying a uniform translation to each map element (dividers, boundaries, crosswalks); as for the predicted online local maps, in actual scenarios, the online local maps are generated by vehicles passing through the same area, and their data is stored on cloud servers for other vehicles to retrieve in the future. For datasets collected by a single vehicle such as nuScenes, Argoverse 2, and OpenLane-V2, certain road sections are captured through multiple drives. Therefore, we use each baseline model to infer the entire dataset, thereby generating online local prior maps at different trips or different times in the same trip, simulating this actual process. Due to the different occlusion conditions of different trips, the online local prior map contains more abundant supplementary information for the current driving scene. So we first search the historical prediction of different trips, and then search the historical prediction of the past time in the same trip as the local priors.
@@
PriorDrive 71.8 81.4 74.1 75.8
     w/o UVE 67.3 74.0 67.8 69.7
     w/o Pretrain 69.4 76.5 69.6 71.5
     w/o SD Prior 69.5 77.1 72.2 72.9
     w/o Online Local 68.4 73.9 69.5 70.6
     w/o HD Prior 67.3 72.0 68.6 69.3
Table 4: The ablation of each component of PriorDrive.
Main Results
@@
We evaluate our method across different model, evaluation metrics in Table 1. Our PriorDrive achieved a 7.2 mIoU improvement over baseline HDMapNet, which demonstrates that our vector prior remains effective even for rasterized models and metrics. In single-variable comparisons with prior/temporal-based methods employing MapTR as baseline, PriorDrive has demonstrated the most substantial improvement of 16 mAP. We further benchmarked our approach against the mainstream MapTRv2 baseline, which demonstrated an average performance gain greater than those achieved on HDMapNet, demonstrating that our HPQuery achieves more comprehensive and fine-grained exploitation of prior information. Furthermore, PriorDrive outperforms existing prior/temporal-based SOTA methods that adopt the MapTRv2 baseline, demonstrating the effectiveness of PriorDrive in fully utilizing complementary information from different prior maps. Figure 5 shows that using the prior maps can effectively recover the occluded elements, as well as more accurate mapping results. Overall, our method demonstrated consistent performance improvements across all baselines, highlighting its generalizability and potential applicability to other frameworks.
Figure 5: Comparison of qualitative results between those with and without unified prior maps.
@@
In Table 2, we present the results of 3 dim maps on Argoverse 2. PriorDrive improves performance by 8.1 mAP and outperforms the previous prior/temporal-based SOTA methods, demonstrating excellent generalization performance.
@@
We compared PriorDrive on OpenLane-V2 when using ResNet-50 for training 24 epochs in Table 3. Compared to TopoLogic baseline, our method achieves better topological reasoning accuracy. And our overall score OLS reached 46.2 (v1.0.0) and 48.5 (v2.1.0), outperforming both Topologic and previous prior-based approaches. Overall, PriorDrive is a plug-and-play solution that can be widely applied to other vector tasks.
Ablation Study
we conducted ablation experiments on nuScenes based on MapTRv2 unless otherwise specified. More experiments are included in supplement.
Ablation Study of PriorDrive.
The ablation study in Table 4 investigates the contribution of each component. Specifically, replacing the UVE with separate MLPs for each prior degrades performance by 6.1 mAP, demonstrating UVE’s effectiveness in mitigating the inconsistency and weak alignment among different priors. indicating it can also alleviate these issues and enhance generalization. The performance degradation observed when ablating any single prior highlights that they contain complementary information, which PriorDrive is the first to leverage simultaneously.
Method $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
StreamMapNet 29.6 30.1 41.9 33.9
HRMapNet 36.9 30.3 44.0 37.1
MapTracker 45.9 30.0 45.1 40.3
PriorDrive (Ours) 44.7 35.5 47.3 42.5
Table 5: Comparisons on the new nuScenes split data sets proposed in StreamMapNet.
Comparison on New Split.
The original split of nuScenes exhibits geographical overlap between training and validation sets. Therefore, we adopt new split proposed by StreamMapNet in Table 5, which training and validation data are separated in locations. PriorDrive achieves 42.5 mAP, which outperforms the previous SOTA methods.
Method Range $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
[…]
StreamMapNet 37.2 42.3 30.2 36.6
HRMapNet 43.1 47.7 34.9 41.9
PriorDrive (Ours) 45.0 49.5 37.3 43.9
[…]
StreamMapNet 22.4 31.3 16.0 23.3
HRMapNet 29.4 34.2 19.7 27.8
PriorDrive (Ours) 28.8 36.3 21.6 28.9
Table 6: Performance comparison of large perception ranges on nuScenes.
Comparison of Large Perception Ranges.
Existing online mapping methods are constrained in prediction range due to the low resolution of distant areas in the captured images which leads to a significant decline in performance. However, PriorDrive leverages various prior maps to enhance accuracy in long-range predictions in Table 6, demonstrating an outstanding foresight capability.
Conclusion
In this paper, we introduced PriorDrive, a novel framework that effectively and uniformly leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction. Through comprehensive experiments, we demonstrated that UVE, combined with proposed pre-training strategies, significantly improves performance of state-of-the-art models in various vector tasks. PriorDrive not only addresses the challenges associated with online HD map construction in complex environments but also offers a scalable solution that continually improves accuracy over time. The iterative use of historical prediction maps as priors leads to progressively refined map outputs, providing latest road information for autonomous driving.
Supplementary Material
Appendix A More Experiments
Here, we present a detailed examination of additional ablation studies on MapTRv2 unless otherwise specified.
M N Dim AP
@@ Appendix A More Experiments
w/o UVE 67.3 74.0 67.8 69.7
4 4 64 71.3 80.0 74.2 75.2
2 2 128 71.0 80.2 73.5 74.9
2 2 64 71.8 81.4 74.1 75.8
1 2 64 70.2 79.1 73.2 74.2
2 1 64 70.4 78.5 71.9 73.6
1 1 32 68.1 76.3 69.6 71.3
Table 7: Ablation experiments of UVE structure.
Structure of UVE.
Table 7 presents the results of ablation experiments on the UVE structure. As the UVE model parameters are progressively augmented, the mAP increases to 75.8, while the model maintains a lightweight architecture. However, performance saturates with further parameter increases, constrained by the limited volume and diversity of the vector data.
Method AP
@@
w/o Pretrain 69.4 76.5 69.6 71.5
Original - - 69.9 77.1 70.2 72.4
Noise 3m 20% - 69.8 78.0 71.7 73.2
Noise 1m 10% - 71.0 79.2 72.7 74.3
Noise 1m 10% 5% 72.0 80.1 73.5 75.2
Mask 10% 5% 71.4 79.8 73.5 74.9
Mask & Noise 1m 10% 5% 71.8 81.4 74.1 75.8
Table 8: Ablation of the pre-training strategies. “Seg” and “Pt” represent segment-level and point-level, respectively.
The Effect of Pre-training.
To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies in Table 8. When we fed the original vector map directly to UVE without any noise, the simple task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.9 mAP improvement. We randomly add segment-level or point-level Gaussian noise to map elements. Among them, 10% segment-level noise and 5% point-level noise increase 3.7 mAP the most. This demonstrates that strategy incorporating both segment-level and point-level noise are more effective at learning the prior information. We also experimented with masking instead of noise, which resulted in a 3.4 mAP improvement. Finally, we conducted pre-training by simultaneously applying gaussian noise and mask, achieving the best result of 75.8 mAP.
@@
w/o online local None 68.4 73.9 69.5 70.6
search_range=5m One 70.6 78.6 71.7 73.6
pretrain=true Two 70.6 79.2 72.4 74.1
integration=replace Multiple 71.8 81.4 74.1 75.8
Table 9: Ablation study on the number of online local prior maps.
Impact of the Number of Prior Maps.
In history, different vehicles passed through the same area. Due to the varying degrees of obstruction, these vehicles would predict different online local maps, thus being able to be used as prior maps. Also, the online local maps predicted by the same vehicle a few seconds ago also have reference value, so they can also be used as prior maps. Therefore, we conducted an ablation study to assess the effect of the number of online local prior maps (see Table 9). The results indicate that using a single online local prior map led to a 3.0 mAP improvement, demonstrating its effectiveness in enhancing current perception. Incorporating two prior maps resulted in a more substantial improvement of 3.5 mAP, reflecting the value of richer prior information. Due to the varying number of available online local prior maps in different regions, and because of the hyperparameter settings of UVE that can encode vector points, the number of multiple online local prior maps is generally 3 to 5. Multiple prior maps effectively improved 5.2 mAP, indicating that our UVE can adaptively alleviate potential inconsistencies across different online local maps and more prior maps can provide more comprehensive necessary information.
@@
w/o online local None 68.4 73.9 69.5 70.6
prior_num=2 5 m 71.8 81.4 74.1 75.8
pretrain=true 10 m 70.2 79.4 72.7 74.1
integration=replace 15 m 69.0 76.7 70.8 72.2
@@
Effect of different search ranges for prior maps.
We search for available online local prior maps based on the distance between the geographical location when the online local prior map was generated and the current geographical location of the vehicle. So we performed ablation experiments to explore the impact of search ranges for prior maps in Table 10. The results show that as the search range is narrowed, the effect gradually improves. Although a smaller search range may reduce the number of online local prior maps, it will be more accurate in terms of geographical overlap. In such a situation of slight inconsistency and misalignment, our PriorDrive can effectively alleviate the problem.
@@
w/o online local None 68.4 73.9 69.5 70.6
prior_num=2 Replace 71.5 79.6 73.1 74.7
pretrain=true Add 71.8 80.2 73.6 75.2
search_range=5m Concat 71.8 81.4 74.1 75.8
@@
To evaluate the versatility of our approach, we tested various integration methods in Table 11. Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 4.1 and 4.6, respectively, but these methods may impact the performance of the original query. In contrast, the concatenation method achieved a more substantial improvement of 5.2 mAP. This approach effectively preserves the original query’s integrity while also leveraging the instance-level and point-level prior features learned by UVE, making it the most effective integration method.
@@
Table 12: Detailed comparison of our method with MapTRv2.
Inference Speed, Memory, Model Size, and Runtime.
The results in Table 12 measured on a single RTX A6000 show that our method maintains similar FPS, GPU memory usage, and parameter count compared to MapTRv2, while improving 14.3 mAP. Our UVE accounts for only 15.3% of the total pipeline, highlighting the efficiency and practical value of our approach across various application scenarios.
Appendix B More Qualitative Visualization
We visualize map construction results of PriorDrive under various weather conditions on nuScenes in Figure 6. PriorDrive maintains stable and impressive results. Moreover, in Figure 7, Figure 8 and Figure 9, we show the result comparison between MapTRv2 and the proposed PriorDrive on the nuScenes. Finally, we also present the visualizations of Argoverse 2 in Figure 10 and OpenLane-V2 in Figure 11. Our results have more complete map representations, more accurate shape and point positions of map elements.
Figure 6: Our method maintains stable and robust HD map construction quality in complex and variable driving scenarios on nuScenes.
Figure 7: Qualitative result comparison under the weather condition of cloudy on nuScenes.
Figure 8: Qualitative result comparison under the weather condition of rainy on nuScenes.
Figure 9: Qualitative result comparison under the night condition on nuScenes.
Figure 10: Qualitative results on Argoverse2 val set. Since we generate a 3D vectorized HD map, the predictions can be rendered precisely on the surrounding view images.
Figure 11: Qualitative results on OpenLane-V2. The red arrows represent the predicted lane topology.